First Community Bank pays an APR of 6.22% simple interest on its savings account, whereas First Mutual Bank pays an APR of 5.38% compounded semiannually. If you made a deposit of $14,500 in each bank today, which of the bank accounts would be a better investment over a five year investment period ?
Answer #1

First Community Bank :

Simple interest=Principal*Interest rate*time period

=14500*6.22%*5=$4509.5

Future value=Simple interest+Principal

=14500+4509.5=$19009.5

First Mutual Bank:

We use the formula:
A=P(1+r/2)^2n
where
A=future value
P=present value
r=rate of interest
n=time period.

A=14500*(1+0.0538/2)^(2*5)

=14500*1.30401185

=$18908.17(Approx)

Hence better investment would be First Community Bank having higher future value.